# Parametric tuning of dynamical phase transitions in ultracold reactions

### v1

@@ III Quasi-adiabatic second-order phase transition
for a general value of $r>-g$. In terms of the parameter $\kappa=r/g$, this corresponds to the re-scaling $g^{2}\rightarrow g^{2}(1+\kappa)$. The smaller sweep rates, $\beta/(g^{2}+rg)<1/N$, are not captured by the formula (19) because this regime corresponds to the onset of the truly adiabatic evolution, with exponentially suppressed excitations. For experimentally relevant values, $N>10^{4}$, this truly adiabatic regime cannot be reached, so we leave it without further discussions.
The change of the adiabatic invariant is interpreted in terms of the density of the nonadiabatic excitations: $n_{ex}=(N-\langle n\rangle)/N=\Delta I/N$, which in turn is related to the deviation of the average number of the formed molecules from its maximal value. Our analysis so far has been restricted to the quasi-adiabatic regime. A comparison with numerical results is shown in Fig. 2. The agreement for the interval $1/\log(N)\gg(g^{2}+rg)/\beta\gg 1/N$ is indeed found. However, beyond this interval the deviations from our formulas are strong. Note also that these deviations persist to smaller $\beta$ when $\kappa$ approaches the value $\kappa=-1$, indicating the breakdown of our analysis for $\kappa\leq-1$.
In summary, our theory predicts the robustness of the 2nd order dynamical phase transition for repulsive and moderate attractive interactions ($\kappa>-1$). Its signature in experiments can be the scaling of the density of the excitations (unformed molecules) for different sweep rates in the quasi-adiabatic regime:
$n_{ex}\sim\beta\log\beta,$
@@
IV First-order phase transition
The energy (10) has no finite global minimum for $r<-g$, so we must include the next order term, $\propto n^{3}$ in the effective Hamiltonian (10). After the transformations (11) and (12), we find
$H_{\rm eff}^{\text{I}}(Q,P,s):=-\beta s\>Q^{2}+\left(\frac{r+g}{N}\right)Q^{4}+\left(\frac{g}{4N^{2}}\right)Q^{6}+\frac{g}{2}\frac{P^{2}}{2}.$ (21)
Disregarding the higher order terms in $P$ and $Q$ is justified when the main nonadiabatic effects occur for $1\ll Q^{2}\ll N$. This requires that
@@ IV First-order phase transition
By approaching the time moment $s=0$ the minimum at $Q=0$ becomes initially a false vacuum, and for $s>0$ it becomes an unstable local maximum. The steady point at $Q=0$ can then be perturbed by any quantum fluctuation, so at $s=0$ the system has to fall towards one of the remaining minima, which at this moment are given by
@@
with corresponding energy $V(Q_{\pm})=-64|\delta r|^{3}N/(27g^{2})$. The system can choose to be in one of the two minima with equal probability. At this point, let us consider that the system escapes towards $Q_{+}$.
@@
We plot the dependence of $n_{\text{ex}}^{\text{nad}}$ on $\delta r/g$ in the nonadiabatic regime in Fig. 3. Our numerical Fig. 3 was obtained by looking at the changes of $n_{ex}(\delta r)$ at smallest possible but finite $\beta$. Hence, only the slope of the line but not the constant offset from the axes origin could be compared to Eq. (24).
@@
Figure 3: (a) Numerically obtained dependence of the density $n_{\text{ex}}$ of the excitations on $\beta_{\rm eff}$ for different values of the coupling $r$ (at $N=100$). The curves correspond to the optimal fit for the exponent $\mu$ as predicted by (26). (b) Asymptotic dependence of the quasi-adiabatic saturation value of $n_{\text{ex}}$ on the ratio $\delta r^{2}/g^{2}$ in the adiabatic limit ($N=500$). The slope of the linear fit is ${\sim}0.97$, which corresponds well with the analytical expression in (24). (c) Asymptotic dependence of the $\beta$-dependent part of $n_{\text{ex}}$ on $\delta r/g$ for the non-adiabatic case at different values of sweeping rate $\beta_{\text{eff}}$, where $\beta_{\text{eff}}=(\beta\log_{e}(N))/g^{2}$ (decreasing from $1.5$ to $0.6$, top to bottom), the slopes of which are all the same (${\sim}0.22$) as expected ($N=200$).
In general, for a finite sweeping rate $\beta$, we also have a Kibble-Zurek scaling of the number of excitations:
@@
where the exponent $\mu$ depends on $\kappa$. In the Methods section, we estimated this exponent analytically for $r=-g$ ($\kappa=-1$) to be $2/3$, which is different from the exponent $1$ of the second-order phase transition (see Fig. 3 for numerical proof). For $\kappa<-1$, we have numerically, together with (24):
@@
V Transient dynamics and prethermalization
A characteristic dynamical feature accompanying the phase transition at time $t=0$ is the oscillation of atomic and molecular populations which become rapidly damped with time. Such transient oscillations of the number of molecules are usually observed in the nonadiabatic regime but become suppressed in the adiabatic limit for a typical second-order phase transition. However, we found that the transient state in the case of a first-order phase transition remains persistent and highly nonclassical.
This is manifested as a complex series of collapse and revival of the number of molecules as a function of time (see Fig. 4). This pattern is particularly prominent in the adiabatic limit. We note that such patterns have been encountered previously in the time-independent versions of the Jaynes-Cummings model [24], Bose-Hubbard model [25], a model of nonlinear directional couplers [26], and even in a generalized version of the model in (1) but at other conditions [27].
In our model, in the case of the second-order transition, these collapse-revival oscillations are absent (see Fig. 4). The quasi-adiabatic regime then produces a sharply-peaked molecular number distribution. In contrast, the number distribution for the first-order transition is very broad, which indicates a phase squeezing. As discussed for related models in [25, 28], such collapse-revival patterns are a purely quantum-mechanical phenomenon; the semiclassical Hamiltonian in (8) does not predict these features.
@@ V Transient dynamics and prethermalization
Figure 4: All the plots in this figure have been generated for $N=200$. (a) Plot of $\langle n_{\text{mol}}\rangle$ (scaled by $N$) vs time for $\kappa=-3$, i.e. for quasi-adiabatic first-order transition. Inset: Dependence of $\langle n_{\text{mol}}\rangle$ with $t$ in the long-time limit. (b) Coherent projection of the molecular wavefunction (see Methods) at the times marked in red dots in (a) (left-right-inset in (a), corresponds to left-right-top-bottom in (b)). (c) Plots of the number distribution of the molecules for the times marked in (a) (from left to right, followed by inset). The grey plot corresponds to the average asymptotic number distribution in the limit $t\rightarrow\infty$. (d) Same as in (a), but for the second-order phase transition at $\kappa=0$. (e) Same as in (b), but with $\kappa=0$. (f) Same as in (c), but with $\kappa=0$. (g) Plot of the number of the revivals in a given time window for different values of $\kappa$ in the first-order phase transition (same $g$, adiabatic limit), fit to a linear curve (dashed line). (h) Time period of a particular revival as a function of $\kappa$ fit to $T_{\text{rev}}\sim 1/|\kappa|$ (dashed line).
To further understand the characteristics of this collapse-revival oscillations, we generated projections of the time-dependent molecular wavefunction onto the basis of coherent states, as shown in Fig. 4. It reveals that the collapse and revival patterns constitute an oscillation of the number distribution $P(n_{\text{mol}})$ about the asymptotic limit (marked in gray in Fig. 4). We also found that the number $N_{\text{rev}}$ of revivals per unit time increases linearly with the strength of nonlinearity ($|\kappa|$) for the first-order transition. The frequency of the collapse-revival patterns decreases gradually following the sweep, and in general saturates to a constant value resulting in a punctuated series of revivals that survive for a very long time. The time-period $T_{\text{rev}}$ of any particular collapse and revival pair also decreases with $|\kappa|$ as
$T_{\text{rev}}\sim\frac{1}{|\kappa|}.$ (28)
We plot the numerically found dependence of $N_{\text{rev}}$ and $T_{\text{rev}}$ on $\kappa$ in Fig. 4. The sustained oscillation pattern and the dependence of the frequency and time period on the interaction strength are reminiscent of the prethermalized states [29] that were numerically observed [30, 31] following a sudden quench in bosonic/fermionic Hubbard models. A main aspect of such nonequilibrium states is the existence of long-time memory of the initial conditions as in our model.
One of the main consequences of this long-term memory and inherently nonequilibrium dynamics of the prethermalized state is the generation of a quantum cat state in the $t\rightarrow\infty$ limit of the forward sweep (see Fig. 4 showing the snapshots for time-evolution in the Glauber basis). In the limit $t\rightarrow+\infty$, after the adiabatic first-order transition, the dynamics eventually freezes in a superposition of two macroscopically distinguishable states (shown as the bottom-right panel in Fig. 4). Each of these states in the Glauber basis is localized, and therefore is similar to a macroscopic Bose-Einstein condensate with a well defined phase. However, the phases of the condensates in the cat state are different by $\pi$. Thus, even the prethermalized state of the nonadiabatic excitations is strongly nonclassical.
VI Discussion
We have explored the role of molecular interactions in mediating dynamic phase transitions in ultracold chemistry with an extension of the Tavis-Cummings model. The standard integrable driven Tavis-Cummings model has made several predictions which should be observable in experiments that associate ultra-cold atoms into molecules by a stimulated passage through the Feshbach resonance. We tested the robustness of these predictions against additional molecular interactions that break the model’s integrability. For moderate interactions we found that, as in the integrable case, the system passes through the second-order phase transition. In the quasi-adiabatic regime, this leads to a nearly linear scaling, $n_{ex}\propto\beta\log_{e}\beta$, of the number of the unformed molecules on the sweep rate $\beta$.
Above a certain critical interaction strength, the order of the phase transition changes from the second to the first, with drastic changes in the reaction dynamics. The main characteristic of the crossover to the first-order phase transition is that even in the adiabatic limit the density of the excitations remains finite. Accompanying this is a transient dynamics that manifests as pronounced oscillations in the molecular population. Asymptotically, the system reaches a prethermalized state (Fig. 4), which is a strongly nonclassical superposition of two condensate-like states with a phase difference of $\pi$ between them, a feature which is absent in the typical case of the second-order phase transition.
Such dynamical features, which are usually attributed to the strongly nonadiabatic regime (e.g., following a sudden quench, as in [30, 31, 32, 29, 33]), are observed in our model (2), during the quasi-adiabatic transition through a first-order critical point. This has a potential impact on the endeavours in creating non-classical states such as the macroscopic cat states and squeezed states in bosonic systems, which are of utmost importance in quantum sensing and metrology [34]. In the past, several experiments have targeted the creation of such non-classical states [35, 36, 37], some of which used a Kerr-type nonlinearity to stabilize these states. Given that the size of an atomic cloud is of the order of $N>10^{4}$, the passage through the Feshbach resonance with a first-order phase transition presents an attractive opportunity to create practically desired strongly nonclassical macroscopic states.
Apart from applications in quantum sensing, experiments in ultracold chemistry also serve to demonstrate the universality of the dynamics around quantum phase transitions [38]. For instance, there is considerable evidence (both theoretical and observational) that phase transitions occurred during the early evolution of our universe [39]. Our study reveals that purely quantum correlations play a considerable role in such processes, and that the nature of the quantum critical point is manifested in the scaling of the excitation density. The main experimental challenge in studying such correlations will be in tuning the interaction strength and obtaining the desired degree of quantum control. With major strides of experimental development in ultracold chemistry in the last decade, we expect this to be possible in the near future.
VII Methods
VII.1 Numerical simulations
The time-dependent Schrödinger equation that describes the time evolution of the Hamiltonian (2) was propagated using the Trotter-factorized discretization of the unitary propagator. Since (2) is also sparse (as is evident from (5)), the time complexity of the cost of the propagation scales almost linearly with $N$.
VII.2 Husimi-Q projection
The coherent Husimi Q projection of the wavefunction in the molecule number basis (Fig. 4) was obtained as:

### v2

@@ II.2 Quasi-adiabatic second-order phase transition
for a general value of $r>-g$. In terms of the parameter $\kappa=r/g$, this corresponds to the rescaling $g^{2}\rightarrow g^{2}(1+\kappa)$. The smaller sweep rates, $\beta/(g^{2}+rg)<1/N$, are not captured by the formula (21) because this regime corresponds to the onset of the truly adiabatic evolution, with exponentially suppressed excitations. For experimentally relevant values, $N>10^{4}$, this truly adiabatic regime cannot be reached, so we leave it without further discussion.
The change of the adiabatic invariant is interpreted in terms of the density of the nonadiabatic excitations: $n_{ex}=(N-\langle n\rangle)/N=\Delta I/N$, which in turn is related to the deviation of the average number of the formed molecules from its maximal value. Our analysis so far has been restricted to the quasi-adiabatic regime. A comparison with numerical results is shown in Fig. 2. We indeed find agreement with the theory for the interval $1/\ln(N)\gg(g^{2}+rg)/\beta\gg 1/N$. However, beyond this interval, the deviations from our formulas are strong. Note also that these deviations persist to smaller $\beta$ when $\kappa$ approaches the value $\kappa=-1$, indicating the breakdown of our analysis for $\kappa\leq-1$.
In summary, our theory predicts the robustness of the dynamics during a passage through a second-order quantum phase transition for repulsive and moderate attractive interactions ($\kappa>-1$). Its signature in experiments can be the scaling of the density of excitations (unformed molecules) for different sweep rates in the quasi-adiabatic regime:
$n_{ex}\sim\beta\ln\beta,$
@@
II.3 First-order phase transition
The energy (12) has no finite global minimum for $r<-g$, so we must include the next order term, $\propto n^{3}$ in the Hamiltonian (5). After the transformations (13) and (14) in the classical limit, we find
$\begin{split}H_{\rm eff}^{\text{I}}(Q,P,s):=-\beta s\>Q^{2}+\left(\frac{r+g}{N}\right)Q^{4}\\ +\left(\frac{g}{4N^{2}}\right)Q^{6}+\frac{g}{2}\frac{P^{2}}{2}\end{split}$ (23)
Disregarding the higher-order terms in $P$ and $Q$ is justified when the main nonadiabatic effects occur for $1\ll Q^{2}\ll N$. This requires that
@@ II.3 First-order phase transition
By approaching the time moment $s=0$, the minimum at $Q=0$ becomes initially a false vacuum, and for $s>0$ it becomes an unstable local maximum. The steady point at $Q=0$ can then be perturbed by any quantum fluctuation, so at $s=0$ the system has to fall towards one of the remaining minima, which, at this moment are given by
@@
with the corresponding energy $V(Q_{\pm})=-64|\delta r|^{3}N/(27g^{2})$. The system can choose to be in one of the two minima with equal probability. At this point, let us consider that the system escapes towards $Q_{+}$.
@@
We plot the dependence of $n_{\text{ex}}^{\text{nad}}$ on $\delta r/g$ in the nonadiabatic regime in Fig. 3. The numerical results in Fig. 3 were obtained by looking at the changes of $n_{ex}(\delta r)$ at the smallest possible but finite $\beta$. Hence, only the slope of the line (and not the constant offset from the origin) could be compared to Eq. (26).
@@
Figure 3: Dependence of the density of nonadiabatic excitations on the sweeping rate and interaction strengths (a) Numerically obtained dependence of the density $n_{\text{ex}}$ of the excitations on $\beta_{\rm eff}$ for different values of the coupling $r$ (at $N=100$). The curves correspond to the optimal fit for the exponent $\mu$ as predicted by (29). (b) Asymptotic dependence of the quasi-adiabatic saturation value of $n_{\text{ex}}$ on the ratio $\delta r^{2}/g^{2}$ in the adiabatic limit ($N=500$). The slope of the linear fit is ${\sim}0.97$, which corresponds well with the analytical expression in (26). (c) Asymptotic dependence of the $\beta$-dependent part of $n_{\text{ex}}$ on $\delta r/g$ for the non-adiabatic case at different values of sweeping rate $\beta_{\text{eff}}$, where $\beta_{\text{eff}}=(\beta\ln N)/g^{2}$ (decreasing from $1.5$ to $0.6$, top to bottom), the slopes of which are all the same (${\sim}0.22$) as expected ($N=200$).
In general, for a finite sweeping rate $\beta$ at $\delta r=0$, we also have a Kibble-Zurek scaling of the number of excitations:
@@
where the exponent $\mu$ depends on $\kappa$. In section IV.4, we estimated this exponent analytically for $r=-g$ ($\kappa=-1$) to be $2/3$, which is different from the exponent $1$ of the second-order phase transition (see Fig. 3 for numerical proof). For $\kappa<-1$, we have numerically, together with (26):
@@
II.4 Transient dynamics and prethermalization
While the type of the phase transition can be understood using the scaling of the number of excitations obtained from the asymptotic number distribution of molecules, the transient dynamics, in the lead up to the number distribution as $t\rightarrow+\infty$, also exhibits features specific to the underlying phase transition. One characteristic feature accompanying the passage through the quantum critical point at time $t=0$ is the oscillation of atomic and molecular populations which become rapidly damped with time. Such transient oscillations of the number of molecules are usually observed in the nonadiabatic regime with a finite sweep rate $\beta>0$. In fact, the frequency dependence of such nonadiabatic oscillations between atomic and molecular populations was used in [8] to infer the role of bosonic enhancement of the ultracold reaction. These oscillations are expected to vanish in the adiabatic limit $\beta\rightarrow 0$ as per the mean-field theory detailed in previous sections, as the dynamics is supposed to ‘equilibrate’ to one of the minima of the mean-field potentials in (12) and (23) in the $t\rightarrow\infty$ limit.
In order to understand this transient quantum dynamics and compare it with mean-field theoretic predictions, we plot the numerically obtained instantaneous average number of molecules computed as:
$n_{\text{avg}}(t)=\frac{1}{N}\sum_{n=0}^{N}n\;|\langle n|\psi_{\text{mol}}(t)\rangle|^{2}$ (30)
for the quasi-adiabatic passage through a first-order (Fig. 4) and second-order (Fig. 4) quantum critical point. Here |$\ket{\psi_{\text{mol}}}(t)$ is computed as |$\ket{\psi_{\text{mol}}(t)}=e^{i\hat{H}t}\ket{\psi}_{n=0}$ where |$\ket{\psi}_{n=0}$ corresponds to the initial state with no molecules and $\hat{H}$ is the Hamiltonian in (2). Numerically, this is implemented using a Trotter-factorized discretization of the unitary propagator $e^{i\hat{H}t}$. (Since (2) is also sparse, as is evident from (5), the time complexity of the cost of the propagation scales almost linearly with $N$.)
For the case of a second-order phase transition, Fig. 4 (plotted for $\kappa=0$) shows that $n_{\text{avg}}(t)$ increases monotonically with time to its saturation value of unity indicating a complete conversion of atoms to molecules as predicted by mean-field theory in the adiabatic limit. Fig. 4 also shows that the molecule number distribution $P(n_{\text{mol}})$ computed as:
$P({n_{\text{mol}}})=|\langle n_{\text{mol}}|\psi\rangle|^{2}$ (31)
(where |$\ket{\psi}$ is the molecular wavefunction at various instances of time) becomes increasingly sharply peaked as $t\rightarrow\infty$.
However, we found that the transient state in the case of a first-order phase transition remains persistent and highly nonclassical. This is manifested as a complex series of collapse and revival of the number of molecules as a function of time (see Fig. 4). This beating pattern is particularly prominent in the adiabatic limit and is distinct from the non-adiabatic oscillations reported in [8]. We note that such patterns have been encountered previously in the time-independent versions of the Jaynes-Cummings model [32], Bose-Hubbard model [33], a model of nonlinear directional couplers [34], and even in a generalized version of the model in (1) but for other starting conditions [35].
In stark contrast to the dynamics around a second-order critical point, the number distribution for the first-order transition is very broad even as $t\rightarrow\infty$, which indicates a phase squeezing. As discussed for related models in [33, 36], the existence of such collapse-revival patterns is a purely quantum-mechanical phenomenon; the semiclassical Hamiltonian in (10) does not predict these features.
@@ II.4 Transient dynamics and prethermalization
Figure 4: Characteristics of the quantum dynamics during the passage through a critical point (All the plots in this figure have been generated for $N=200$). (a) Plot of $\langle n_{\text{mol}}\rangle$ (scaled by $N$) vs time for $\kappa=-3$, i.e. for quasi-adiabatic first-order transition. Inset: Dependence of $\langle n_{\text{mol}}\rangle$ with $t$ in the long-time limit. (b) Coherent projection of the molecular wavefunction (see Methods) at the times marked in red dots in (a) (left-right-inset in (a), corresponds to left-right-top-bottom in (b)). (c) Plots of the number distribution of the molecules for the times marked in (a) (from left to right, followed by inset). The grey plot corresponds to the average asymptotic number distribution in the limit $t\rightarrow\infty$. (d) Same as in (a), but for the second-order phase transition at $\kappa=0$. (e) Same as in (b), but with $\kappa=0$. (f) Same as in (c), but with $\kappa=0$. (g) Plot of the number of the revivals in a given time window for different values of $\kappa$ in the first-order phase transition (same $g$, adiabatic limit), fit to a linear curve (dashed line). (h) Time period of a particular revival as a function of $\kappa$ fit to $T_{\text{rev}}\sim 1/|\kappa|$ (dashed line).
To further understand the aspects of quantum dynamics that lead to this collapse-revival oscillation, we generated Husimi Q projections of the time-dependent molecular wavefunction onto the basis of coherent states, as shown in Fig. 4 (for second-order transition) and Fig. 4 (for first-order transition). The coherent Husimi Q projections of the wavefunction in the molecule number basis is obtained as: